OpenStudy (anonymous):

Suppose the function g(x) is used to model the height y, of a ball x seconds after the ball is thrown up in the air. Suppose further that a ball reaches its maximum height of 7.5 feet in 5 seconds, and then hits the ground 3 seconds later. What would be an appropriate domain for g(x)?

OpenStudy (anonymous):

I thinking its is 0 ≤ x ≤ 8 because domain is x-values and range is the y-values. Not 100% sure, my domain n range skills are rusty.

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

The max height occurs at the vertex

jimthompson5910 (jim_thompson5910):

The vertex lies on the axis of symmetry

jimthompson5910 (jim_thompson5910):

So this means that there is a vertical line that cuts the parabola (which opens downward) in 2 exact mirroring halves at x = 5

jimthompson5910 (jim_thompson5910):

So if it takes 3 seconds (from x = 5) to hit the ground, which is the x-axis, then it should take 3 seconds for the ball to launch into the air and hit the peak So the domain is actually 2 <= x <= 8 because 2 and 8 are both 3 seconds away from x = 5 seconds

OpenStudy (anonymous):

im sorry i had to step away from the computer

OpenStudy (anonymous):

thanks your answer makes alot more sense and i get it thanks alot!

jimthompson5910 (jim_thompson5910):

that's great, glad you do

OpenStudy (anonymous):

wait i have multiple choice and thats not a choice

jimthompson5910 (jim_thompson5910):

what are your choices?

OpenStudy (anonymous):

there is no 2 <= x <= 8

OpenStudy (anonymous):

they are 3< or equal to x<or equal to 5 -5<or equal to x<or equal to 5 0<or equal to x<or equal to 8 0<or equal to x< or equal to 7.5

jimthompson5910 (jim_thompson5910):

well it can't be 0 to 8 because the midpoint of this interval is 4 (because 8+0 = 8 and you divide this in half to get 4)

jimthompson5910 (jim_thompson5910):

this is an odd problem

OpenStudy (anonymous):

I know im really stuck

jimthompson5910 (jim_thompson5910):

however, if you use x = 0 to x = 8 and you discount x = 0 to x = 2, then I guess it could work (in an odd way)

jimthompson5910 (jim_thompson5910):

but there is no physical way to have a ball reach its peak in 5 seconds, come back down in 3 seconds and have it start on the ground at 0 seconds...it's just not possible

OpenStudy (anonymous):

so your guess would be 0<=x<=8

jimthompson5910 (jim_thompson5910):

yeah that's if you ignore the fact that x = 0 to x = 2 (excluding x = 2 itself) don't work

jimthompson5910 (jim_thompson5910):

which is why this is an odd problem

OpenStudy (anonymous):

could you help with one more!

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

So I can get that one right!

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

A biologist took count of the nuber of migrating waterfowl at a particular lake and recounted the lakes population of waterfowl on each of the next six weeks.

OpenStudy (anonymous):

week: 0 1 2 3 4 5 6 Population: 635 644 719 860 1,067 1,340 1,679

OpenStudy (anonymous):

this is the table

OpenStudy (anonymous):

Find a quadratic function that models the data as a function of x, the number of weeks. Use the model to estimate the number of waterfowl at the lake on week 8.

jimthompson5910 (jim_thompson5910):

are you familiar with polynomial interpolation?

OpenStudy (anonymous):

no whats that!

jimthompson5910 (jim_thompson5910):

it's what is needed to solve this problem polynomial interpolation is the method to find the polynomial (in this case, the quadratic) that goes through every point

OpenStudy (anonymous):

no i dont

jimthompson5910 (jim_thompson5910):

hmm seems odd how you need to use something you haven't learned yet

OpenStudy (anonymous):

could you explain it

jimthompson5910 (jim_thompson5910):

sure I can do that

OpenStudy (anonymous):

these are my answers p(x)=33x^2-24x+635:2,555 waterfowl p(x)=38x^2+24x+585: 2615 p(x)=33x^2-24x+635:2,084 p(x)=38x^2+24x+585: 3,209

jimthompson5910 (jim_thompson5910):

well since you have the answer choices, we can do it a much simpler way if you want

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

ok, for each answer choice, the number at the very end is going to be ignored for now, so erase those numbers for each answer choice to get p(x)=33x^2-24x+635 p(x)=38x^2+24x+585 p(x)=33x^2-24x+635 p(x)=38x^2+24x+585

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Now look at the last numbers of each choice They are: 635, 585, 635, 585

jimthompson5910 (jim_thompson5910):

These are the y - intercepts of each answer choice

OpenStudy (anonymous):

so how do i know which one to pick

jimthompson5910 (jim_thompson5910):

remember that the y-intercepts are the points (0, b) where b is some number

jimthompson5910 (jim_thompson5910):

because we have the point (0, 635), this means that we can rule out choices B and D because they have the y-intercept of (0, 585)

jimthompson5910 (jim_thompson5910):

So we now have these two answer choices A) p(x)=33x^2-24x+635 C) p(x)=33x^2-24x+635

jimthompson5910 (jim_thompson5910):

they are the same polynomial luckily, so this means that the answer to the first part is p(x)=33x^2-24x+635

jimthompson5910 (jim_thompson5910):

to answer the second part, you plug in x = 8 and evaluate/simplify

OpenStudy (anonymous):

so its A because it equals 2555

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

you make it seem so easy :)

jimthompson5910 (jim_thompson5910):

thanks lol

jimthompson5910 (jim_thompson5910):

the other way would have done the trick too, but it was much longer/harder, so i figured i'd go this way

OpenStudy (anonymous):

Can you check my answer for this one. there is a table x -2 0 4 f(x) 0 -6 78

OpenStudy (anonymous):

i have to write equation indard form of parabola that models the values of the table I got y=5x^2+4x-6

OpenStudy (anonymous):

in standard form sorry!

OpenStudy (anonymous):

is that right

jimthompson5910 (jim_thompson5910):

unfortunately no, here's why plug in x = -2 to get... 5x^2+4x-6 5(-2)^2+4(-2)-6 5(4)+4(-2)-6 20 - 8 - 6 12 - 6 6 So 5x^2+4x-6 = 6 when x = -2 BUT we want 5x^2+4x-6 to equal 0 when x = -2 So it can't be the answer because of this reason.

OpenStudy (anonymous):

ohh so how could i do it

OpenStudy (anonymous):

my other answers choices are y=6x^2+5x-4 y=-4x^2-5x+6 Y=4x^2+5x-6

OpenStudy (anonymous):

is there a way i can work backwards or something

jimthompson5910 (jim_thompson5910):

well you know that one point is (0, -6)

jimthompson5910 (jim_thompson5910):

so the y-intercept is (0, -6) which means that the last number must be a -6

OpenStudy (anonymous):

so in this case it has to be the last one

jimthompson5910 (jim_thompson5910):

so that leaves the last choice y=4x^2+5x-6 sure enough, plugging in x = -2 gives you y=4x^2+5x-6 y=4(-2)^2+5(-2)-6 y=0 which is what we want (so far so good) and plugging in x = 4 gives you y=4x^2+5x-6 y=4(4)^2+5(4)-6 y=78 which is also what we want So all 3 points lie on this parabola

OpenStudy (anonymous):

I cant explain how grateful i am! thanks so much!!!

OpenStudy (anonymous):

You really were a big help!!

jimthompson5910 (jim_thompson5910):

yw

jimthompson5910 (jim_thompson5910):

I'm glad that I was

jimthompson5910 (jim_thompson5910):

hmm interesting, thx for letting me know

OpenStudy (anonymous):

@lala2 to answer the confusion of the first question you asked and @jim_thompson5910 the reason the solution is 0<x<8 is because the x axis is the second and y is the feet high. So he would have thrown the ball at 0 seconds not 2 seconds.